### v2

# Avoiding strict saddle points of nonconvex regularized problems

Abstract
In this paper, we consider a class of non-convex and non-smooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms. These problems, moreover, are probably non-lipschitz around sparse solutions. We propose an damped iterative reweighted $\ell_{1}$ algorithm to solve these problems. The algorithm is guaranteed to converge only to local minimizers when randomly initialized. By deriving a second-order sufficient optimal condition under the non-degenerate assumption, we show that the strict saddle property is generic on these sparse optimization problems.
keywords:
strict saddle, iterative reweighted $\ell_{1}$ algorithm, nonconvex regularization, center stable manifold theorem
1 Introduction
In this paper, we consider the following nonsmooth regularization problem,
$\min\limits_{x\in\mathbb{R}^{n}}F(x):=f(x)+\lambda\psi(x),$ (P)
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$. We assume the objective $F$ has bounded level sets, the data fitting function $f$ is twice continuously differentiable and possibly nonconvex, and $r$ satisfies the following properties.
Assumption 1.
$r$ is continuous on $[0,+\infty)$ with $r(0)=0$, $C^{2}$-smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz around 0.
The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance. This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4]. Many widely used nonconvex regularization such as the smoothly clipped absolute deviation penalty (SCAD) [5] or the minimax concave penalty (MCP) [6], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability. We provide the explicit form of several commonly used nonconvex regularization in Table 1. It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) non-Lipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as $\ell_{p}$ quasi-norm (LPN).
Numerous studies focus on gradient-based optimization methods that are guaranteed to find first-order stationary points. However, the nonconvexity of $F$ can lead to multiple local minima and saddle points, which can cause algorithms to converge to stationary points rather than the (local) optimum. Additionally, the presence of saddle points can significantly slow down the convergence of optimization algorithms, especially gradient-based methods. At a saddle point, the gradient provides misleading information about the descent directions. As a result, avoiding saddle points can help optimization algorithms more effectively converge to better solutions. Therefore, it becomes a prevalent topic, particularly in high-dimensional optimization problems, and is also the central focus of this paper.
In smooth optimization, many methods have been developed and analyzed to escape from saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19]. In smooth settings, it is widely assumed that the strict saddle property holds, meaning all saddle points possess at least one strictly negative eigenvalue. This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases and has been shown to be generally satisfied for smooth problems [20, 21].
In nonsmooth optimization, avoiding saddle points poses a greater challenge. At stationary point, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate the avoidance of saddle points. Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization. An example of this complexity is demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [22]. Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point. Additionally, several proximal methods, including proximal point method, proximal gradient method and proximal linear method [23], as well as a more efficient algorithm incorporating perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings. In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$-weakly convex meaning $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. They assume an active-manifold around the limit point and then define the strict saddle points as those of the $f$ restricted to the manifold. The weakly $\rho$-convexity for $r$ plays a crucial role in constructing C2 differentiable Moreau envelope with the same critical points and strict saddle points in proximal methods.
As for the regularized problem (P), several first-order methods have been proposed. [25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1. This method is subsequently extended to address group $\ell_{p,q}$ norm regularized problems in [26]. Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions. It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold true for the $\ell_{p}$ regularized problem. Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularizers. Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [28] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems [29]. Subsequent work in [30] established its global convergence and demonstrated its locally stable support property [31]. Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [32] and its extrapolated version was proposed in [33]. While these first-order methods primarily focus on identifying first-order critical points, none of them guarantee avoidance of saddle points.
1.1 The main results
In this paper, our main purpose is to develop an iterative reweighted $\ell_{1}$ algorithm that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity. The additional assumption for $r$ needed in our work to guarantee that the existence of an active manifold around an active manifold as assumed by [23].
Assumption 2.
For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior).
The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) is widely applicable to these non-convex regularization problems. In contrast to the proximal gradient algorithm, ${\rm IRL}_{1}$ solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator. This method offers a favorable local property [31], known as the model identification. At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0. In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{J}$ as follows after finite number of iterations.
Definition 1.
(Model Subspace) The model subspace defined by $\mathcal{J}\subset\{1,\ldots,n\}$ is
$\mathcal{M}_{\mathcal{J}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{J},u_{j}=0\}.$
This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{J}}$. If we think the iteration of ${\rm IRL}_{1}$ as a mapping that assigns to every point another point. This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) on $\mathcal{M}_{\mathcal{J}}$, even though the underlying optimization problem is nonsmooth in the entire space.
We can summarize our contribution as follows:
  1. (i)
    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{J}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{J}}$.
  2. (ii)
    We develop damped IRL1 algorithms individually for the two types of regularizers, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning the next iterate is the combination of a traditional IRL1 iterate and its previous iterate. The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ becomes $\infty$ at each iteration.
  3. (iii)
    At the tail end of the iterations, we show this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{J}$ on the subspace $\mathcal{M}_{\mathcal{J}}$ and decrease the rest components to 0 at a linear speed. We then apply the center stable manifold theorem and prove that the algorithm can avoid converge to strict saddle points.
1.2 Outline
[…]
1.3 Notation
[…]
${\rm sign}(t)=\left\{\begin{aligned} &1,\ &t>0,\\ &-1,\ &t<0,\\ &0,\ &t=0.\end{aligned}\right.$ (1)
$DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$. Given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}\subset\{1,\ldots,n\}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants. If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$. Given an index set $\mathcal{I}$, denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively. Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$. Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the vector of thresholds $w$ is defined $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$. In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.
Regularizer $\psi(x)$ $r^{\prime}(|x_{i}|)$
EXP [34] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $pe^{-p|x_{i}|}$
LOG [35] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$
FRA [36] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$
LPN [7] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $p|x_{i}|^{p-1}$
SCAD [5] $r(|x_{i}|)=\lambda\left\{\begin{array}[]{ll}|x_{i}|&{\rm if}\ |x_{i}|\leq\lambda\\ \frac{a\lambda-|x_{i}|}{a-1}&{\rm if}\ \lambda<|x_{i}|\leq a\lambda\\ \frac{(a+1)\lambda^{2}}{2}&{\rm if}\ |x_{i}|>a\lambda\end{array}\right.$ $\lambda\{\mathbb{I}(|x_{i}|\leq\lambda)+\frac{(a\lambda-|x_{i}|)_{+}}{(a-1)\lambda}\mathbb{I}(|x_{i}|>\lambda)\}$
MCP [6] $r(|x_{i}|)=\left\{\begin{array}[]{ll}\lambda|x_{i}|-\frac{\beta^{2}}{2a}&{\rm if}\ |x_{i}|\leq a\lambda\\ \frac{a\lambda^{2}}{2}&{\rm otherwise}.\end{array}\right.$ $(a\lambda-|x_{i}|)_{+}/a$
Table 1: Concrete examples of nonconvex regularization
2 Preliminaries
2.1 Optimality condition
In this subsection, we introduce the first and the second-order optimality condition of problem (P). First, let us recall the definition of subgradient and the first-order optimality condition.
Definition 2.
[…]
  1. (1)
    $v$ is a regular subgradient of $f$ at $\bar{x}$, written $v\in\hat{\partial}f(\bar{x})$ if
    $\partial f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
  2. (2)
    $v$ is a (general) subgradient of $f$ at $\bar{x}$, written $v\in\partial f(\bar{x})$, if there are sequences $x^{v}\rightarrow\bar{x}$ and $v^{v}\in\hat{\partial}f(x^{v})$ with $x^{v}\rightarrow v$.
Let $s(\cdot)=r\circ|\cdot|$. It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$.
The subgradients of $s$ are given in the following proposition.
Proposition 1.
It holds that $\partial F(x)=\hat{\partial}F(x)=\nabla f(x)+\lambda\partial\psi(x)$ where
$\partial\psi(x)=\hat{\partial}\psi(x)=\partial\sum\limits_{i=1}^{n}s(x_{i})=\partial s(x_{1})\times...\times\partial s(x_{n})$ (3)
and
$\partial s(x_{i})=\hat{\partial}s(x_{i})=\begin{cases}\{\text{sign}(x_{i})r^{\prime}(|x_{i}|)\},&i\in\mathcal{I}(x),\\ [-r^{\prime}(|0^{+}|),+r^{\prime}(|0^{+}|)],&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)<\infty,\\ \mathbb{R},&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)=\infty.\end{cases}$ (4)
Proof.
We only need calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
$\hat{s}(x):=\begin{cases}r(|x|)&\text{ if }x\geq 0,\\ r^{\prime}(|0^{+}|)x&\text{ otherwise}.\end{cases}$
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [37, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, meaning $r^{\prime}(|0^{+}|),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [37, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
If $r^{\prime}(|0^{+}|)=\infty$, we have $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{r(|x^{*}|)-r(0)-v(x-0)}{|x-0|}\geq 0$ for all $v\in\mathbb{R}$. Therefore, we have ^$\mathbb{R}\subseteq\hat{\partial}s(0)\subseteq\partial s(0)\subseteq\mathbb{R}$. Hence, ^$\hat{\partial}s(0)=\partial s(0)=\mathbb{R}$.
Overall, we have shown that (4) is true. It then follows from [37, Proposition 10.5] that (3) is true. ∎
We can appeal to the classical rule of Fermat [37, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a critical point of (P), then
$0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$ (5)
This condition, combined with Assumption 2 and the explicit form of the subgradients, yields the following results.
Proposition 2.
If $x^{*}$ is a local minimum for (P), then the following holds
$\nabla_{i}f(x^{*})=-\text{sign}(x_{i}^{*})\lambda r^{\prime}(|x^{*}_{i}|),$ $i\in\mathcal{I}^{*},$ (6)
$\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$ (7)
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the critical point.
In the following, we demonstrate that the second-order optimality condition can be characterized by the nonzero components.
Theorem 3.
  1. (i)
    (necessary) If $x^{*}$ is a local minimum of $F$, then
    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
    or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))\geq 0$.
  2. (ii)
    (sufficient) A critical point $x^{*}$ of $F$ is a strict local minimum if
    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
    or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>0$.
Proof.
(i) is proved in [38, Theorem 3.3]. We only prove (ii).
Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of non-zero $\{u^{k}\}$ satisfying
$F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$ (10)
Consider $F_{|\mathcal{I}^{*}}$, which is locally smooth around $x^{*}_{\mathcal{I}^{*}}$ and $x^{*}_{\mathcal{I}^{*}}$ is stationary for $F_{|\mathcal{I}^{*}}$. It follows from $u^{k}=u_{\mathcal{I}^{*}}^{k}+u_{\mathcal{J}^{*}}^{k}$ that
$F(x^{*}+u^{k})-F(x^{*})$
$=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+f(x^{*}+u_{\mathcal{I}^{*}}^{k})$
$-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the critical point of (P) and (8) holds. Therefore for sufficiently large $k$,
$F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$ (11)
Moreover, by the Lagrangian mean value theorem, there exists sufficiently small $\epsilon^{k}\in(0,\min_{i}\{|u_{i}^{k}|\})$ such that
$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$
The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$; consequently,
$\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{1}\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{2}.$
This, combined with (11), yields
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$ (12)
where $\eta^{k}:=\frac{f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})}{\|u_{\mathcal{J}^{*}}^{k}\|}$. By the definition of gradient,
$\lim_{u^{k}\to 0}\frac{f(x^{*}+u_{\mathcal{J}^{*}}^{k})-f(x^{*})-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\lim_{u^{k}\to 0}\eta^{k}-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=0,$
implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\|\nabla f(x^{*})\|_{\infty}$. By (7), $\limsup\limits_{k\to\infty}\eta^{k}=\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$. Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$. It follows from (12) that for sufficiently large $k$,
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|>0,$
contradicting (10). ∎
2.2 Strict Saddle Point
In this section, we define the strict saddle point of the problem (P). In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction. For our case, Theorem 3 implies a saddle point can be defined using the nonzeros.
Definition 3.
A critical point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathcal{M}_{\mathcal{I}^{*}}\setminus\{0\}$. If all critical points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points. To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the critical point on which the function is partly smooth. Our definition also belongs to this kind, as $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold containing $x^{*}$.
2.3 Center Stable Manifold Theorem
As many other works in smooth setting like [40], we interpret our algorithm as the fixed point iteration of a nonlinear map $T:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$. Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
Definition 4.
(Unstable fixed point) Let $x^{*},\epsilon^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. $x^{*},\epsilon^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*},\epsilon^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points. The CSM theorem establishes a favorable property for the fixed point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
Theorem 4.
(The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$. Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one. There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$, and $\cap_{k=0}^{\infty}T^{-k}(B)\subset W_{loc}^{cs}$.
The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. Such result was established in [40, Theorem 2] for a global diffeomorphism map, under which the preimage of a measure zero set also has measure zero. The following corollary in [23] shows the same conclusion for lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous).
Corollary 1.
[23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible. Then, the set of initial points attracted by such fixed points
$W:=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$
has zero Lebesgue measure.
3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
The ${\rm IRL}_{1}$ algorithm is an effective method to solve problem (P). At each iteration, the ${\rm IRL}_{1}$ algorithm solves the following weighted $\ell_{1}$ regularized subproblem,
$x^{k+1}=S(x^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i})|y_{i}|,$
where $w(x_{i})=r^{\prime}(|x_{i}|)$. To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a critical point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally non-invertible. To see this, one can rewrite the map $S$ using the soft-threshold operator with a changing threshold as
$S(x^{k})=\mathbb{S}_{\lambda w^{k}}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})).$
This algorithm may quickly converges to a local minimizers due to the rapid growth of the weight values, especially for the case when $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered. To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
$\min\limits_{\theta\in\mathbb{R}^{2n}}F(\theta):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}^{2}).$ (13)
with $\theta:=[x;\epsilon]$ for shorthand. Here we also treat $\epsilon$ as another variable to optimize, so that the algorithm updates both $x$ and $\epsilon$ by solving the following subproblem
$x^{k+1}=S(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w_{i}^{k}|y_{i}|,$
with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and setting $\epsilon^{k+1}=\mu\epsilon^{k}$ with $\mu\in(0,1)$.
The subproblem solution can be given explicitly using the solft-thresholding operator. This update can be written as
$S(x^{k})=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (14)
or explicitly,
$S_{i}(x^{k})=\left\{\begin{aligned} &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ &0,&\text{otherwise }.\end{aligned}\right.$ (15)
The map $S$ is not a bijection, as it maps multiple values to 00. To address this issue, we need to revise this algorithm to ensure it satisfy these conditions. Motivated by the technique used in [23], we modified the algorithm with following damping iteration,
$x^{k+1}=(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k}),$
where $\alpha\in(0,1)$ is the damping parameter. Intuitively, if $S_{i}(x^{k},\epsilon^{k}_{i})=0$, the damping operator maps $x^{k}_{i}$ to $(1-\alpha)x^{k}_{i}$, resulting in a bijective map. We can use the iterate $(x;\epsilon)\in\mathbb{R}^{2n}$ and equivalently write the update of $(x^{k};\epsilon^{k})$ as
$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k})\\ \mu\epsilon^{k}\end{bmatrix}$ (16)
with $\epsilon_{0}>0$, $\mu\in(0,1)$ and $\alpha\in(0,1)$.
This algorithm, hereinafter nicknamed DIRL1, is presented in Algorithm 10. If the algorithm does not involve the vector of $\epsilon$, we can simply treat this case as setting $\epsilon^{0}=0$ in Algorithm 10 and the iteration simply reverts to
$x^{k+1}=T(x^{k})=(1-\alpha)x^{k}+\alpha S(x^{k}).$
We have the following result to connect the optimality of (P) and the fixed point of our iteration.
Proposition 5.
$x^{*}$ is a critical point of (P) if and only if $x^{*}=S(x^{*},0)$. In addition, $x^{*}$ is a critical point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
Proof.
Suppose $x^{*}$ is a critical point. It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}(x^{*},0)=x_{i}^{*}$. This result also holds true by the same argument if $x_{i}^{*}<0$.
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}(x^{*},0)=0=x_{i}^{*}$. Overall, we have shown that $x^{*}=S(x^{*},0)$ in all cases.
Conversely, suppose $x^{*}=S(x^{*},0)$. If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$. Same argument yields $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$. Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
If $x_{i}^{*}=0$, it indicates that $|\frac{1}{\beta}\nabla_{i}f(x^{*})|=|x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})|\leq\tfrac{\lambda}{\beta}w(0,0)=\tfrac{\lambda}{\beta}r^{\prime}(|0^{+}|)$. Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$. Overall, $x^{*}$ satisfies (5) and is a critical point of (P). By Assumption 2, it satisfies (6) and (7).
The rest part of the statement of this proposition is obviously true by noticing that $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$ implies that $\epsilon^{*}=0$. ∎
1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
2:Initialize: set $k=0$.
3:repeat
4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})$.
5:     Compute new iterate:
6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}\{\nabla f(x^{k})^{T}(y-x^{k})+\frac{\beta}{2}\|y-x^{k}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})|y_{i}|\}$.
7:     $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$.
8:     $\epsilon^{k+1}=\mu\epsilon^{k}$.
9:     Set $k=k+1$.
10:until Convergence
Algorithm 1 Damped Iteratively Reweighted $\ell_{1}$ Algorithm (${\rm DIRL}_{1}$)
4 Convergence analysis
The analysis focuses solely on the algorithm involving the update of $\epsilon$. It is noteworthy the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
Assumption 3.
Given the initial point $(x^{0}$, $\epsilon^{0})$ and $F^{0}=F(x^{0},\epsilon^{0})$, the level set $\mathcal{L}(F^{0}):=\{(x,\epsilon)\mid F(x,\epsilon)\leq F^{0}\}$ is bounded.
This assumption means $f$ is Lipschitz continuous and Lipschitz differentiable on $\mathcal{L}(F^{0})$ with constant $L_{f}$, i.e., $\forall x,y\in\mathbb{R}^{n}$ (we use the same constant for $f$ and $\nabla f$ for simplicity).
$\|f(x)-f(y)\|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (17)
We first show the monotonicity of $F(x,\epsilon)$ along a direction. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
$\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}),$
$\Delta G(x^{k+1};x^{k},\epsilon^{k})=$ $\ G(x^{k};x^{k},\epsilon^{k})-G(x^{k+1};x^{k},\epsilon^{k}).$
Proposition 6.
Suppose Assumption 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
Moreover, $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$ and there exists $C>0$ such that
$\|x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})\|_{\infty}\leq C,$ (18)
for any $k\in\mathbb{N}$.
Proof.
Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we can imply that for $i=1,...,n$,
$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})+r^{\prime}(|x_{i}^{k}|+(\epsilon^{k}_{i})^{2})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
Summing up the above inequality over $i$ yields
$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (19)
On the other hand, it follows that
$f(x^{k})-f(x^{k+1})\geq\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (20)
Combining (19) and (20), we have
$\Delta F(x^{k+1},\epsilon^{k+1})\geq F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k})$ (21)
$\geq$ $\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)$
$=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is by the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
For each subproblem, $y^{k}$ satisfies the optimal condition
$\nabla f(x^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})=0.$ (22)
with $\xi^{k}\in\partial\|y^{k}\|_{1}$ and $W(x^{k},\epsilon^{k}):={\rm diag}(w(x_{1}^{k},\epsilon_{1}^{k}),...,w(x_{n}^{k},\epsilon_{n}^{k}))$. Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (22), the following holds
$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (23)
$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(\alpha|x_{i}^{k}|-\alpha|y_{i}^{k}|)-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})\xi_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ \alpha[\nabla f(y^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is by $\alpha\in(0,1)$ and the second is by the convexity of ||$\cdot$|| ⋅ |. Combine (21), (23) and get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
Replacing $k$ with $t$ and summing up from $t=0$ to $k-1$, we have
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
It follows that $\{(x^{k},\epsilon^{k})\}\in\mathcal{L}(F^{0})$ since $F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$. By Assumption 3, it is bounded and $\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})>-\infty$. Taking $k\rightarrow\infty$, we have
$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, it follows from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (18) holds for any $k\in\mathbb{N}$. ∎
With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
Theorem 7.
(Global convergence) Suppose Assumption 1 and 3 hold. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\mu^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
Proof.
Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$
Then we consider the two kinds of cases with $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$.
If $r^{\prime}(|0^{+}|)=\infty$, as shown in Proposition 2, we only need to consider $i\in\mathcal{I}^{*}$. Taking the limit on $\mathcal{S}$,
00 0 $=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(y_{i}^{k})+\beta(y_{i}^{k}-x_{i}^{k})|$
$=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})+\frac{\beta}{\alpha}(x_{i}^{k+1}-x_{i}^{k})|$
$\geq\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})|-\frac{\beta}{\alpha}\|x^{k}-x^{k+1}\|$
$=|\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0){\rm sign}(x_{i}^{*})|$
$=|\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|){\rm sign}(x_{i}^{*})|.$
which satisfy (5). Therefore, $x^{*}$ is stationary.
If $r^{\prime}(|0^{+}|)<\infty$, we need also to consider $i\in\mathcal{J}^{*}$. By [41, Proposition 2.1.5] and [37, Theorem 5.19], $\partial r$ is outer semi-continuous and there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
$0=\nabla_{i}f(x^{*})+\lambda w(0,0)\xi^{*}_{i}=\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)\text{sign}(x_{i}^{*}),$
meaning (5) is satisfied. Therefore, $x^{*}$ is stationary. ∎
5 Avoidance of strict saddle point
DIRL1 has the property of so-called “active manifold identification”, which belongs to the research line of active-manifold identification in nonsmooth optimization [42, 43]. It means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consists of the nonzeros at a limit point.
Theorem 8.
(Model Identification and Local Smoothness) Under Assumption 1-3, for any critical point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
  1. (i)
    For any point $(x,\epsilon)\in B$, $S(x,\epsilon)\in(x^{*}+\mathcal{M}_{\mathcal{I}^{*}})$ and $\text{sign}(S(x,\epsilon))=\text{sign}(x)=\text{sign}(x^{*})$.
  2. (ii)
    Map $S$ is $C^{1}$-smooth in $B$ and $S(x)=x-\frac{1}{\beta}g(\theta)$ with
    $g_{i}(\theta)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
  3. (iii)
    Map $T$ is $C^{1}$-smooth in $B$.
Proof.
(i) Since $S$ is a continuous map and $x^{*}=S(x^{*},\epsilon^{*})$, it holds that
$\text{sign}(x_{i})=\text{sign}(x^{*}_{i})\text{\ and\ }\text{sign}(S_{i}(x,\epsilon))=\text{sign}(S_{i}(x^{*},0))=\text{sign}(x^{*}_{i}),\quad i\in\mathcal{I}^{*}$
for any $(x,\epsilon)$ sufficiently close to $(x^{*},\epsilon^{*})$. It remains to show that for any $i\in\mathcal{J}^{*}$, $S_{i}(x,\epsilon)=0$ for any $(x,\epsilon)\in B$. By Assumption 2, $|x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})|=|\tfrac{1}{\beta}\nabla_{i}f(x^{*})|<\tfrac{1}{\beta}\lambda w(0,0)$. Therefore, there exists a neighborhood $B$ such that $|x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)|<\tfrac{1}{\beta}\lambda w(x_{i},\epsilon_{i})$, $\forall(x,\epsilon)\in B$, which is equivalent to $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))<0$ and $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))>0$, meaning $S_{i}(x,\epsilon)=0$.
(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agree with following constrain problem
$S(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (24)
$=$ $\ \arg\min\limits_{x\in\mathbb{R}^{n}}\{g(x^{k},\epsilon^{k})^{T}x+\frac{\beta}{2}\|x-x^{k}\|^{2}\},$
where each component of $g:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ is defined by
$g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
Therefore, $S$ is $C^{1}$-smooth in $B$
(iii) It follows from (24) that the iteration $T$ is $C^{1}$-smooth in $B$ for any for any $(x,\epsilon)\in B$. ∎
The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a critical point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$. Specifically, it can be expressed as
$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &\mu\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (25)
around $(x^{*},\epsilon^{*})$. We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima. To apply the CSM theorem, we have to show the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
Theorem 9.
(Unstable Fixed Point) Suppose $x^{*}$ is a critical point of (P). $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$.
Proof.
First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 8. It holds that
$\frac{\partial g}{\partial\epsilon_{i}}(x,\epsilon)=\begin{cases}0&\text{ \ if}i\in\mathcal{J}^{*},\\ \lambda r^{\prime\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*})2\epsilon&\text{ \ if }i\in\mathcal{I}^{*}.\end{cases}$
Let $D_{\epsilon}g(x,\epsilon)$ be the Jacobian matrix of $g(\cdot)$ with respect to $\epsilon$. It follows that $D_{\epsilon}g(x^{*},\epsilon^{*})=\bm{0}$ by the fact that $\epsilon_{i}^{*}=0$ and $r^{\prime\prime}(|x_{i}^{*}|)<+\infty$, $i\in\mathcal{I}^{*}$. On the other hand, $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$. Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
$DT^{*}(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &\mu I\end{bmatrix}=\begin{bmatrix}I-\frac{\alpha}{\beta}D_{x}g(x^{*},\epsilon^{*})\ &\bm{0}\\ \bm{0}\ &\mu I\end{bmatrix}$
$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&\mu I\end{bmatrix}$
Now we are ready to prove the equivalence between the unstable fixed point of $T$ and the strict saddle point of (P). Recall Definition 4 and Definition 3 and notice $\mu\in(0,1)$ and $\alpha\in(0,1)$. It suffices to show that $I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has an eigenvalue greater than 1 if and only if $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has a negative eigenvalue. This equivalence is obviously true since
$\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0\iff\lambda_{\max}(I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>1.$
To apply Corollary 1, we next to show $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$. This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
Proposition 10.
Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ defined in Proposition 6, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. $T$ is a lipeomorphism when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
Proof.
If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i}^{2})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
If $r^{\prime}(|0+|)=\infty$, by Proposition 6, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, then it must be true that $S_{i}(x,\epsilon)=0$. Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
$r^{\prime}(|x_{i}|)\leq C\text{ \ and \ }|x_{i}|\geq\underline{x}:=(r^{\prime})^{-1}(C)>0$ (26)
by Assumption 3.
This implies that we can equivalently write $S$ as
$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
with
$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i}^{2})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}^{2}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
Therefore, $u(\cdot)=\min(r^{\prime}(\underline{x}),r^{\prime}(\cdot))$ is also Lipschitz continuous in $\mathcal{L}(F^{0})$. Let $L_{r}$ be the Lipschitz constant, meaning for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$
$\|u(\hat{x},\hat{\epsilon})-u(\tilde{x},\tilde{\epsilon})\|\leq L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$ (27)
A natural choice is $L_{r}=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$ and $L_{r}=r^{\prime\prime}(|\underline{x}|)$ if $r^{\prime}(|0^{+}|)=\infty$.
Next we prove $S$ is Lipschitz continuous. It follows that
$S(\hat{x},\hat{\epsilon})-S(\tilde{x},\tilde{\epsilon})=$ $\ \mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))$
$\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\|\hat{u}-\tilde{u}\|$
$\leq$ $\ \|\hat{x}-\tilde{x}\|+\tfrac{1}{\beta}\|\nabla f(\hat{x})-\nabla f(\tilde{x}))\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f})\|\hat{x}-\tilde{x}\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
where the first inequality is by Lemma A.2 and the second is by (17) and (27). Therefore, $\|\alpha(\hat{x}-S(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$. Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is a lipeomorphism. ∎
Now we are ready to apply Corollary 1 to derive the following global escape theorem.
Theorem 11.
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3. Let the condition of Proposition 10 holds. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure. of (P).
Proof.
It follows directly from Theorem 9 that map $T$ is local $C^{1}$-smooth around any critical points, and all critical points $x^{*}$ are unstable fixed points. By applying Proposition 10, $T$ is a lipeomorphism. By Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure. ∎
Appendix
Lemma A.1.
Given $a,b,c\in\mathbb{R}^{n}_{+}$ with $a_{i}\leq b_{i}+c_{i}\quad\text{for all }i$. It holds that $\|a\|\leq\|b\|+\|c\|$.
Proof.
It follows that $a_{i}^{2}\leq b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2}$, implying $\sum_{i=1}^{n}a_{i}^{2}\leq\sum_{i=1}^{n}(b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2})$, or equivalently, $\|a\|^{2}\leq\|b\|^{2}+2b^{T}c+\|c\|^{2}$. By the Cauchy-Schwarz inequality, $2\sum_{i=1}^{n}b_{i}c_{i}\leq 2\|b\|\|c\|$. Therefore, $\|a\|^{2}\leq(\|b\|+\|c\|)^{2}$, completing the proof. ∎
Lemma A.2.
Given $\hat{w},\tilde{w}\in\mathbb{R}^{n}_{+}$ and $\hat{z},\tilde{z}\in\mathbb{R}^{n}$, it holds that
$\|\mathbb{S}_{\hat{w}}(\hat{z})-\mathbb{S}_{\tilde{w}}(\tilde{z})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$
Proof.
(a) If $|\hat{z}_{i}|\leq\hat{w}_{i}$ and $|\tilde{y}_{i}|\leq\tilde{w}_{i}$, then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=0$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$. In this case, $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=|0-0|=0\leq|\hat{z}_{i}-\tilde{z}_{i}|$.
(b) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|\leq\tilde{w}_{i}$ (or vice versa), then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$. In this case, $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=||\hat{z}_{i}|-\hat{w}_{i}|$. It follows that
$|\hat{z}_{i}-\tilde{z}_{i}|\geq$ $\ |\hat{z}_{i}|-|\tilde{z}_{i}|\geq|\hat{z}_{i}|-\tilde{w}_{i}$
$=$ $\ ||\hat{z}_{i}|-\hat{w}_{i}+\lambda(\hat{w}_{i}-\tilde{w}_{i})/\beta|$
$\geq$ $\ |\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|-|\lambda(\hat{w}_{i}-\tilde{w}_{i})/\beta|,$
implying
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
(c) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|>\tilde{w}_{i}$, then
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})-\text{sign}(\tilde{z}_{i})(|\tilde{z}_{i}|-\tilde{w}_{i})|.$
If $\text{sign}(\hat{z}_{i})=\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|(|\hat{z}_{i}|-\hat{w}_{i})-(|\tilde{z}_{i}|-\tilde{w}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|$. If $\text{sign}(\hat{z}_{i})=-\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=(|\hat{z}_{i}|-\hat{w}_{i})+(|\tilde{z}_{i}|-\tilde{w}_{i})\leq|\hat{z}_{i}|+|\tilde{z}_{i}|=|\hat{z}_{i}-\tilde{z}_{i}|.$
In all three cases, we show
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
Therefore, by (A.1),
$\|\mathbb{S}_{\hat{w}}(\hat{w})-\mathbb{S}_{\tilde{w}}(\tilde{w})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$