### v4

Abstract
In this paper, we consider a class of nonconvex and nonsmooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms. We show the second-order optimality conditions only depend on the nonzeros of the stationary points. We propose two damped iterative reweighted algorithms, including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems. For DIRL1, we show the reweighted $\ell_{1}$ subproblem has support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point. For DIRL2, we show the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere. Therefore, the map of DIRL1 and DIRL2 and their inverse are Lipschitz continuous, and the strict saddle points are their unstable fixed points. By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with random initialization when the strictly saddle point property is assumed.
keywords:
strict saddle, iterative reweighted $\ell_{1}$ algorithm, nonconvex regularization, center stable manifold theorem
1 Introduction
In this paper, we consider the following nonsmooth regularization problem,
$\min\limits_{x\in\mathbb{R}^{n}}F(x):=f(x)+\lambda\psi(x),$ (P)
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$. The data fitting function $f$ is differentiable twice continuously and possibly nonconvex, and $r$ satisfies the following properties.
Assumption 1.
$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2 smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz continuous around 0.
The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance. This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4]. Many widely used nonconvex regularization such as the EXP penalty [38] or the LOG penalty [39], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability. We provide the explicit form of several commonly used nonconvex regularizations in Table 1. It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) nonLipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as the $\ell_{p}$ quasi-norm (LPN).
Numerous studies focus on gradient-based optimization methods that are guaranteed to find first-order stationary points. However, the nonconvexity of $F$ can lead to multiple local minima and saddle points, which can cause algorithms to converge to stationary points rather than the (local) optimum. Additionally, the presence of saddle points can significantly slow down the convergence of optimization algorithms, especially gradient-based methods. At a saddle point, the gradient provides misleading information about the descent directions. As a result, avoiding saddle points can help optimization algorithms more effectively converge to better solutions. Therefore, it becomes a prevalent topic, particularly in high-dimensional optimization problems, and is also the central focus of this paper.
In smooth optimization, many methods have been developed and analyzed to escape saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19]. In smooth settings, it is widely assumed that strict saddle property holds, which means that all saddle points have at least one strictly negative eigenvalue. This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases, and it has been shown to be generally satisfied for smooth problems [20, 21].
In nonsmooth optimization, avoiding saddle points poses a greater challenge. At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate avoidance of saddle points. Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization. An example of this complexity is demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [22]. Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point. Additionally, several proximal methods, including the proximal point method, the proximal gradient method, and the proximal linear method [23], as well as a gradient method incorporating random perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings. In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$ weakly convex, which means that $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. They assumed an active manifold around the limit point and then defined the strict saddle points as those of the $f$ restricted to the manifold. The weakly $\rho$-convexity for $r$ plays a crucial role in constructing the C2 differentiable Moreau envelope with the same stationary points and strict saddle points in the proximal methods.
As for the regularized problem (P), several first-order methods have been proposed. [25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1. This method is subsequently extended to address the $\ell_{p,q}$ group norm regularized problems of the group $\ell_{p,q}$ in [26]. Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions. It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold true for the $\ell_{p}$ regularized problem. Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularizers. Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [28] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems [29]. Subsequent work in [30] established its global convergence and demonstrated its locally stable support property [31]. Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [32] and its extrapolated version was proposed in [33]. While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of saddle points.
1.1 The main results
In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity. In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [34, 35, 36, 30, 37, 29, 31, 33]. In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to various nonconvex regularization problems.
The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator. This method offers a favorable local property [31], known as the model/support identification, under the following additional assumption.
Assumption 2.
For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior).
This kind of assumption is also widely made to guarantee that the existence of an active manifold around a stationary point as assumed by [23]. At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0. In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{I}$ as follows after a finite number of iterations.
Definition 1.
(Model Subspace) The model subspace defined by $\mathcal{I}\subset\{1,\ldots,n\}$ is
:=$\mathcal{M}_{\mathcal{I}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{I}^{c},u_{j}=0\}.$
This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{I}}$. This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) in $\mathcal{M}_{\mathcal{I}}$, although the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these nonconvex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given. Compared to ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property. In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero. To show the different behaviors of the two algorithms, we provide the following example.
Example 1.
Consider the $\ell_{1/2}^{1/2}$ regularization problem
$\min_{x\in\mathbb{R}^{2}}F(x)=f(x)+\|x\|_{1/2}^{1/2},\ \text{ with }f(x)=x_{1}^{2}+(x_{2}-5/4)^{2}.$
The global minimizer is $(0,1)$, the local minimizer111(0,0)00(0,0)( 0 , 0 ) is always a dummy local minimizer of $\ell_{p}$ regularization problems is $(0,0)$ and the saddle point minimizer is $(0,\frac{3-2\sqrt{2}}{4})\approx(0,0.043)$ which is also a local maximizer on the $x_{2}$-axis.
(a) Contour of $F$.
(b) Contour of $F$ around $(0,0)$.
Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$. The iterations of the iteratively reweighted algorithms are as follows [29]. We use the ${\rm IRL}_{1}$ algorithm with iteration
$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{2}\frac{1}{(|x_{1}^{k}|+0.1^{k})^{1/2}}|x_{1}|+\tfrac{1}{2}\frac{1}{(|x_{2}^{k}|+0.1^{k})^{1/2}}|x_{2}|$
and the ${\rm IRL}_{2}$ algorithm with iteration
$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{4}\frac{1}{((x_{1}^{k})^{2}+0.1^{k})^{3/4}}x_{1}^{2}+\tfrac{1}{4}\frac{1}{((x_{2}^{k})^{2}+0.1^{k})^{3/4}}x_{2}^{2}.$
The trajectories of the two algorithms are depicted in Figure 2.
(a) Trajectory of ${\rm IRL}_{1}$
(b) Trajectory of ${\rm IRL}_{2}$
Figure 2: Trajectories generated by iteratively reweighted methods. ${\rm IRL}_{1}$ eventually falls into the subspace (the $x_{2}$-axis), while ${\rm IRL}_{2}$ has nonzero elements for all iterates.
The behavior of ${\rm IRL}_{2}$ brings another challenge in analyzing the avoidance of strict saddle point. Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures the algorithm reverts to a smooth algorithm in the active manifold in a local neighbourhood of the stationary point. However, there is no answer how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
We now summarize our contribution as follows:
  1. (i)
    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{I}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{I}}$.
  2. (ii)
    We develop damped IRL1 algorithms individually for the two types of regularizer, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning that the next iterate is the combination of a traditional IRL1 iterate and its previous iterate. The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration. At the tail end of the iterations, we show that this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{I}$ on the subspace $\mathcal{M}_{\mathcal{I}}$ and decrease the rest components to 0 at a linear speed. We then apply the center stable manifold theorem and prove that the algorithm can avoid the convergence to strict saddle points.
  3. (iii)
    We develop damped IRL2 algorithms for nonconvex regularized problems and provide global convergence. Under an additional assumption which is satisfied by most common nonconvex regularizers, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere. We also proved the equivalence between the unstable fixed point of the iteration map and the strict saddle points of the regularization problem. Therefore, the avoidance of strict saddle points is derived. This is the first algorithm shown to have this property without using the active-manifold identification property or the stochastic technique in nonsmooth optimization algorithms.
1.2 Outline
[…]
1.3 Notation
[…]
We denote $\|\cdot\|_{p}$ as the quasi-norm $\ell_{p}$ with $p\in(0,1)$, that is, $\|x\|_{p}=(\sum_{i=1}^{n}|x_{i}|^{p})^{\frac{1}{p}}$. This does not define a proper norm because of the lack of subadditivity. For $x\in\mathbb{R}^{n}$, $x_{i}$ denotes the $i$-th element of $x$. The index set of nonzero and zero components of $x$ is defined as $\mathcal{I}(x)=\{i\mid x_{i}\neq 0\}$, and $\mathcal{J}(x)=\{i\mid x_{i}=0\}$, respectively. Moreover, $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ denotes $\mathcal{I}(x^{*})$ and $\mathcal{J}(x^{*})$ for short. ${\rm diag}(x)$ is the diagonal matrix by placing elements of vector $x$ on the main diagonal. ${\rm sign}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the sign function defined by
$[{\rm sign}(x)]_{i}=\left\{\begin{aligned} &1,\ &x_{i}>0,\\ &-1,\ &x_{i}<0,\\ &0,\ &x_{i}=0.\end{aligned}\right.$ (1)
$DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$. Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}^{*}}\subset\mathbb{R}^{\mathcal{I}^{*}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants. If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}f(x)=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$. Denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively. Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$. Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the vector of thresholds $w$ is defined $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$. In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.
Regularizer $\psi(x)$ $r^{\prime}(|x_{i}|)$ $r^{\prime\prime}(|x_{i}|)$
EXP [38] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $pe^{-p|x_{i}|}$ $-\frac{1}{p^{2}}e^{-\frac{|x_{i}|}{p}}{p}$
LOG [39] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$ $-\frac{1}{(|x_{i}|+p)^{2}}$
FRA [40] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$ $-\frac{2p}{(|x_{i}|+p)^{2}}$
LPN [7] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $p|x_{i}|^{p-1}$ $p(p-1)|x_{i}|^{p-2}$
TAN [41] $\sum\limits_{i=1}^{n}\arctan\frac{|x_{i}|}{p}$ $\frac{p}{(|x_{i}|^{2}+p^{2})^{2}}$ $-\frac{2p|x_{i}|}{(|x_{i}|^{2}+p^{2})^{2}}$
Table 1: Concrete examples of nonconvex regularization
2 Optimality and strict saddle points
2.1 Optimality conditions
In this subsection, we introduce the first and the second-order optimality condition of problem (P). First, let us recall the definition of subgradient and the first-order optimality condition.
Definition 2.
[…]
  1. (1)
    $v$ is a regular subgradient of $f$ at $\bar{x}$, written $v\in\hat{\partial}f(\bar{x})$ if
    ^$\hat{\partial}f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
  2. (2)
    $v$ is a (general) subgradient of $f$ at $\bar{x}$, written $v\in\partial f(\bar{x})$, if there are sequences $x^{v}\rightarrow\bar{x}$ and $v^{v}\in\hat{\partial}f(x^{v})$ with $x^{v}\rightarrow v$.
It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$. The subgradients of $s$ are given in the following proposition.
Proposition 2.
It holds that $\partial F(x)=\hat{\partial}F(x)=\nabla f(x)+\lambda\partial\psi(x)$ where
$\partial\psi(x)=\hat{\partial}\psi(x)=\partial\sum\limits_{i=1}^{n}s(x_{i})=\partial s(x_{1})\times...\times\partial s(x_{n})$ (3)
and
$\partial s(x_{i})=\hat{\partial}s(x_{i})=\begin{cases}\{\text{sign}(x_{i})r^{\prime}(|x_{i}|)\},&i\in\mathcal{I}(x),\\ [-r^{\prime}(|0^{+}|),+r^{\prime}(|0^{+}|)],&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)<\infty,\\ \mathbb{R},&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)=\infty.\end{cases}$ (4)
Proof.
We only need to calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
$\hat{s}(x):=\begin{cases}r(|x|)&\text{ if }x\geq 0,\\ r^{\prime}(|0^{+}|)x&\text{ otherwise}.\end{cases}$
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, which means $r^{\prime}(|0^{+}|)\in\hat{\partial}s(0),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
If $r^{\prime}(|0^{+}|)=\infty$, we have $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{r(|x^{*}|)-r(0)-v(x-0)}{|x-0|}\geq 0$ for all $v\in\mathbb{R}$. Therefore, we have ^$\mathbb{R}\subseteq\hat{\partial}s(0)\subseteq\partial s(0)\subseteq\mathbb{R}$. Hence, ^$\hat{\partial}s(0)=\partial s(0)=\mathbb{R}$.
In general, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
We can appeal to the classical rule of Fermat [42, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a stationary point of (P), then
$0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$ (5)
This condition, combined with Assumption 2 and the explicit form of the subgradients, yields the following results.
Proposition 3.
If $x^{*}$ is a local minimum for (P), then the following holds
$\nabla_{i}f(x^{*})=-\text{sign}(x_{i}^{*})\lambda r^{\prime}(|x^{*}_{i}|),$ $i\in\mathcal{I}^{*},$ (6)
$\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$ (7)
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point that satisfies the necessary optimality conditions (6) and (7) is referred to as the stationary point.
In the following, we demonstrate that the second-order optimality condition can be characterized by the nonzero components.
Theorem 4.
  1. (i)
    (necessary) If $x^{*}$ is a local minimum of $F$, then
    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
    or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))\geq 0$.
  2. (ii)
    (sufficient) A stationary point $x^{*}$ of $F$ is a strict local minimum if
    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
    or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>0$.
Proof.
(i) is proved in [43, Theorem 3.3]. We only prove (ii).
Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of nonzero $\{u^{k}\}$ satisfying
$F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$ (10)
Consider $F_{|\mathcal{I}^{*}}$, which is locally smooth around $x^{*}_{\mathcal{I}^{*}}$ and $x^{*}_{\mathcal{I}^{*}}$ is stationary for $F_{|\mathcal{I}^{*}}$. It follows from $u^{k}=u_{\mathcal{I}^{*}}^{k}+u_{\mathcal{J}^{*}}^{k}$ that
$F(x^{*}+u^{k})-F(x^{*})$ (11)
$=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+f(x^{*}+u_{\mathcal{I}^{*}}^{k})$
$-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
From the classic optimality condition for smooth function and (9), it is easy to see that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$. Therefore for sufficiently large $k$,
$F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$ (12)
Moreover, by the Lagrangian mean value theorem, there exists sufficiently small $\epsilon^{k}\in(0,\min_{i}\{|u_{i}^{k}|\})$ such that
$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$ (13)
The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$, which, together with (13), yields that
$\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{1}\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{2}.$
This, combined with (12) and (11), yields
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$ (14)
where $\eta^{k}:=\frac{f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})}{\|u_{\mathcal{J}^{*}}^{k}\|}$. By the definition of gradient,
$\lim_{u^{k}\to 0}\frac{f(x^{*}+u_{\mathcal{J}^{*}}^{k})-f(x^{*})-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\lim_{u^{k}\to 0}\eta^{k}-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=0,$
implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}\leq\|\nabla f(x^{*})\|_{\infty}$. By (7), $\limsup\limits_{k\to\infty}\eta^{k}\leq\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$. Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$. It follows from (14) that for sufficiently large $k$,
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|>0,$
contradicting (10). ∎
2.2 Strict Saddle Points
In this section, we define the strict saddle point of the problem (P). In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction. For our case, Theorem 4 implies that a saddle point can be defined using nonzeros.
Definition 3.
A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$. If all stationary points of $F(x)$ are strict saddle points or local minimums, $F$ satisfies the strict saddle property.
This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points. To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the stationary point on which the function is partly smooth. Our definition also belongs to this kind, since $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold that contains $x^{*}$.
To show that the strict saddle property is mathematically generic, many works show that this property holds for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2]. Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$. In our case, we also have the same results by considering the following perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
$\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$ (15)
To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7], which is first proved in [46].
Definition 4.
Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, if the derivative at $x$ is singular (Jacobian is not full rank), then $x$ is called a critical point of $f$.
Theorem 5.
(Sard’s Theorem) Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, and let $C$ be the set of stationary points of $f$ in $\mathcal{M}$. Then $f(C)$ has measure zero in $\mathcal{N}$.
The genericity of the strict saddle property is stated below.
Theorem 6.
For a full Lebesgue measure set of perturbations $v\in\mathbb{R}^{n}$, the perturbation function $F_{v}(x):=F(x)-\langle v,x\rangle$ satisfies
$\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d\neq 0,$ (16)
for all stationary points $x$ of $F_{v}$ with $\mathcal{I}=\mathcal{I}(x)$.
Proof.
Let $s\in\{+1,-1,0\}^{n}$ be a vector of sign pattern and
$\mathcal{M}(s)=\{x\in\mathbb{R}^{n}\mid\text{sign}(x)=s\}$
be a manifold. Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$. The Jacobian of $v$ at $x$ is then given by $Dv(x)=\begin{bmatrix}\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)&0\\ 0&I\end{bmatrix}$, and $Dv(x)$ has full rank if and only if $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$ has full rank. By Sard’s Theorem,
$\{v(x)\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n},$
which is equivalent to
$\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}(s)|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}.$
Now we consider all the possible sign pattern vectors $s\in\{+1,-1,0\}^{n}$. There are at most $3^{n}$ different patterns. Therefore,
$\bigcup_{s\in\{+1,-1,0\}^{n}}\mathcal{V}(s)=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathbb{R}^{n}\}$
has measure 0 in $\mathbb{R}^{n}$. ∎
Center Stable Manifold Theorem
As many other works in smooth setting like [47], we interpret our algorithm as the fixed point iteration of a nonlinear map $T$. Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
Definition 5.
(Unstable fixed point) Let $x^{*}$ be a fixed point on a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. $x^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points. The CSM theorem establishes a favorable property for the fixed-point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
Theorem 7.
(The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$. Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one. There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$. In addition, if $T^{k}(x)\in B$ for all $k\geq 0$, then $x\in W_{loc}^{cs}$.
If $\nabla T(0)$ is invertible, then $T$ is locally a diffeomorphism around 0. The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. This requires that the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero. This result was established in [47, Theorem 2] for a global diffeomorphism map and then for a global lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23]. Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown using the fact that $\mathbb{R}^{n}$ is second countable no matter if the number of stationary points is countable (see [23, Corollary 2.12] for a standard proof). We list their results below.
Theorem 8.
[47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$. Let $\mathcal{U}_{T}$ consist of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
Corollary 1.
[23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible. Then, the set of initial points attracted by such fixed points $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim\limits_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
To summarize, the analysis of escaping from strict saddle points under the strict saddle property generally consists of three steps:
  1. (1)
    Interpret the iterative algorithm as a fixed point iteration, and build the equivalence of the unstable fixed points and the strict saddle points.
  2. (2)
    Show that the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at the fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
  3. (3a)
    Show that the fixed-point iteration is globally a diffeomorphism.
  4. (3b)
    (Alternative to (3a)) Show that the fixed-point iteration is globally a lipeomorphism.
3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
The ${\rm IRL}_{1}$ algorithm is an effective method to solve problem (P). At each iteration, the ${\rm IRL}_{1}$ algorithm solves the following weighted $\ell_{1}$ regularized subproblem,
$x^{k+1}=S(x^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i})|y_{i}|,$
where $w(x_{i})=r^{\prime}(|x_{i}|)$.
This algorithm may quickly converge to local minimizers due to the rapid growth of the weight values, especially for the case where $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered. To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
$\min\limits_{x,\epsilon\in\mathbb{R}^{n}}F(x,\epsilon):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}).$ (17)
Here we also treat $\epsilon$ as variables, so that the algorithm updates $x$ and $\epsilon$ by $(x^{k+1},\epsilon^{k+1})=S(x^{k},\epsilon^{k})$
$\begin{cases}x^{k+1}=S^{x}(x^{k},\epsilon^{k})=\arg\min\limits_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum\limits_{i=1}^{n}w_{i}^{k}|y_{i}|,\\ \epsilon^{k+1}=S^{\epsilon}(x^{k},\epsilon^{k})=\mu\epsilon^{k}\end{cases}$
with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and $\mu\in(0,1)$.
The subproblem solution can be given explicitly using the solft-thresholding operator,
$S^{x}(x^{k},\epsilon)=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (18)
or explicitly,
$S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$ (19)
To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally noninvertible. In particular, $S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$ (20)
with $\epsilon_{0}\in\mathbb{R}^{n}_{++}$ and $\alpha\in(0,1)$.
This algorithm, hereinafter nicknamed DIRL1, is presented in Algorithm 10. If the algorithm does not involve the vector of $\epsilon$, we can simply treat this case as setting $\epsilon^{0}=0$ in Algorithm 10 and the iteration simply reverts to
$x^{k+1}=T(x^{k})=(1-\alpha)x^{k}+\alpha S(x^{k}).$
We have the following result to connect the optimality of (P) and the fixed point of our iteration.
Proposition 9.
$x^{*}$ is a stationary point of (P) if and only if $x^{*}=S^{x}(x^{*},0)$. In addition, $x^{*}$ is a stationary point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
Proof.
Suppose $x^{*}$ is a stationary point. It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}^{x}(x^{*},0)=x_{i}^{*}$. This result also holds true by the same argument if $x_{i}^{*}<0$.
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$. In general, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
Conversely, suppose $x^{*}=S^{x}(x^{*},0)$. If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$. Same argument yields $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$. Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
If $x_{i}^{*}=0$, it indicates that $|\frac{1}{\beta}\nabla_{i}f(x^{*})|=|x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})|\leq\tfrac{\lambda}{\beta}w(0,0)=\tfrac{\lambda}{\beta}r^{\prime}(|0^{+}|)$. Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$. In general, $x^{*}$ satisfies (5) and is a stationary point of (P). By Assumption 2, it satisfies (6) and (7).
The rest part of the statement of this proposition is obviously true by noticing that $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$ implies that $\epsilon^{*}=0$. ∎
1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
2:Initialize: set $k=0$.
3:repeat
4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$.
5:     Compute new iterate:
6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}\{\nabla f(x^{k})^{T}(y-x^{k})+\frac{\beta}{2}\|y-x^{k}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})|y_{i}|\}$.
7:     $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$.
8:     $\epsilon^{k+1}=(1-\alpha(1-\mu))\epsilon^{k}$.
9:     Set $k=k+1$.
10:until Convergence
Algorithm 1 Damped Iteratively Reweighted $\ell_{1}$ Algorithm (${\rm DIRL}_{1}$)
3.1 Convergence analysis
It should be noted that the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
Assumption 3.
Given the initial point $(x^{0}$, $\epsilon^{0})$ and $F^{0}=F(x^{0},\epsilon^{0})$, the level set $\mathcal{L}(F^{0}):=\{(x,\epsilon)\mid F(x,\epsilon)\leq F^{0}\}$ is bounded.
This assumption means $f$ is Lipschitz continuous and Lipschitz differentiable on $\mathcal{L}(F^{0})$ with constant $L_{f}$, i.e., $\forall x,y\in\mathbb{R}^{n}$ (we use the same constant for $f$ and $\nabla f$ for simplicity).
$|f(x)-f(y)|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (21)
This assumption also implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (6). Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$. Overall, we can conclude that for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, there exists $\rho$ such that
$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho.$ (22)
We first show the monotonicity of $F(x,\epsilon)$. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
$\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}),$
$\Delta G(x^{k+1};x^{k},\epsilon^{k})=$ $\ G(x^{k};x^{k},\epsilon^{k})-G(x^{k+1};x^{k},\epsilon^{k}).$
Proposition 10.
Suppose Assumptions 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and that the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
Moreover, $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$ and there exists $C>0$ such that
$\|x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})\|_{\infty}\leq C,$ (23)
for any $k\in\mathbb{N}$.
Proof.
Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we have that for $i=1,...,n$,
$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
Summing up the above inequality over $i$ yields
$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+\epsilon_{i}^{k})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+\epsilon_{i}^{k})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (24)
On the other hand, it follows that
$f(x^{k})-f(x^{k+1})\geq\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (25)
Combining (24) and (25), we have
$\Delta F(x^{k+1},\epsilon^{k+1})\geq F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k})$ (26)
$\geq$ $\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)$
$=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is due to the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
For each subproblem, $y^{k}$ satisfies the optimal condition
$\nabla f(x^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})=0.$ (27)
with $\xi^{k}\in\partial\|y^{k}\|_{1}$ and $W(x^{k},\epsilon^{k}):={\rm diag}(w(x_{1}^{k},\epsilon_{1}^{k}),...,w(x_{n}^{k},\epsilon_{n}^{k}))$. Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (27), the following holds
$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (28)
$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(\alpha|x_{i}^{k}|-\alpha|y_{i}^{k}|)-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})\xi_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ \alpha[\nabla f(y^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is by $\alpha\in(0,1)$ and the second is by the convexity of ||$\cdot$|| ⋅ |. Combine (26), (37) and get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
Replacing $k$ with $t$ and summing up from $t=0$ to $k-1$, we have
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
It follows that $\{(x^{k},\epsilon^{k})\}\in\mathcal{L}(F^{0})$ since $F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$. By Assumption 3, it is bounded and $\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})>-\infty$. Taking $k\rightarrow\infty$, we have
$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
Theorem 11.
(Global convergence) Suppose Assumption 1 and 3 hold. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\epsilon^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
Proof.
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$
If $w(x_{i}^{*},0)<+\infty$, it follows from [48, Proposition 2.1.5], [42, Theorem 5.19] and the outer semi-continuity of $\partial r$ that there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
$0=\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)\xi^{*}_{i}$
as $y_{i}^{k}-x_{i}^{k}\overset{\mathcal{S}}{\to}0$, $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$. Therefore, for $w(x_{i}^{*},0)=r^{\prime}(|x_{i}^{*}|)<\infty,i\in\mathcal{I}^{*}$, (6) is satisfied; for $w(x_{i}^{*},0)=r^{\prime}(|0^{+}|)<\infty,i\in\mathcal{J}^{*}$, (7) is satisfied by Assumption 2. Therefore, $x^{*}$ is stationary. ∎
3.2 Avoidance of strict saddle point
The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consisting of nonzeros at a limit point.
Theorem 12.
(Model Identification and Local Smoothness) Under Assumption 1-3, for any stationary point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
  1. (i)
    For any point $(x,\epsilon)\in B$, $S^{x}(x,\epsilon)\in(x^{*}+\mathcal{M}_{\mathcal{I}^{*}})$ and $\text{sign}(S^{x}(x,\epsilon))=\text{sign}(x)=\text{sign}(x^{*})$.
  2. (ii)
    $S$ is $C^{1}$ smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
    $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
  3. (iii)
    Map $T$ is $C^{1}$-smooth in $B$.
Proof.
(i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, we have
$\text{sign}(x_{i})=\text{sign}(x^{*}_{i})\text{\ and\ }\text{sign}(S_{i}(x,\epsilon))=\text{sign}(S_{i}(x^{*},0))=\text{sign}(x^{*}_{i}),\quad i\in\mathcal{I}^{*}$
for any $(x,\epsilon)$ sufficiently close to $(x^{*},\epsilon^{*})$. It remains to show that for any $i\in\mathcal{J}^{*}$, $S_{i}(x,\epsilon)=0$ for any $(x,\epsilon)\in B$. By Assumption 2, $|x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})|=|\tfrac{1}{\beta}\nabla_{i}f(x^{*})|<\tfrac{1}{\beta}\lambda w(0,0)$. Therefore, there exists a neighborhood $B$ such that $|x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)|<\tfrac{1}{\beta}\lambda w(x_{i},\epsilon_{i})$, $\forall(x,\epsilon)\in B$, which is equivalent to $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))<0$ and $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))>0$, meaning $S_{i}(x,\epsilon)=0$.
(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agrees with following constrained problem
$S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (29)
$=$ $\ \arg\min\limits_{x\in\mathbb{R}^{n}}\{g(x^{k},\epsilon^{k})^{T}x+\frac{\beta}{2}\|x-x^{k}\|^{2}\},$
where each component of $g:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ is defined by
$g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
Therefore, $S$ is $C^{1}$ smooth in $B$.
(iii) It follows from (29) that the iteration $T$ is $C^{1}$ smooth in $B$. ∎
The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent method for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$. Specifically, it can be expressed as
$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &(1-\alpha(1-\mu))\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (30)
around $(x^{*},\epsilon^{*})$. We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima. To apply the CSM theorem, we have to show that the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
Theorem 13.
(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P). $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$. Moreover, if $\alpha<\frac{\beta}{\rho}$, $DT(x^{*},\epsilon^{*})$ is invertible.
Proof.
First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 12. It holds that $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$. Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
$DT(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &(1-\alpha(1-\mu))I\end{bmatrix}$
$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}$
Now we are ready to prove the equivalence between the unstable fixed point of $T$ and the strict saddle point of (P). Recall Definition 5 and Definition 3 and notice $\mu\in(0,1)$ and $\alpha\in(0,1)$. It suffices to show that $I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has an eigenvalue greater than 1 if and only if $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has a negative eigenvalue. This equivalence is obviously true since
$\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0\iff\lambda_{\max}(I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>1.$
If $\alpha$ is selected to satisfy $01-\frac{\alpha}{\beta}\lambda_{\max}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>$ 0 and by the strict saddle property $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have 0 eigenvalue, we know $DT(x^{*},\epsilon^{*})$ is invertible. ∎
To apply Corollary 1, we next show that $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$. This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
Proposition 14.
(Lipeomorphism) Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ is defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. Then $T$ is invertible; moreover, $T$ and $T^{-1}$ is Lipschitz continuous (that is, $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
Proof.
If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
If $r^{\prime}(|0+|)=\infty$, by Proposition 10, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, it must be true that $S_{i}(x,\epsilon)=0$. Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
$r^{\prime}(|x_{i}|)\leq C\text{ \ and \ }|x_{i}|\geq\underline{x}:=(r^{\prime})^{-1}(C)>0$ (31)
by Assumption 3. This implies that we can equivalently write $S$ as
$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}(x_{i},\epsilon_{i})/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
with
$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
Therefore, $u(\cdot)=\min(r^{\prime}(\underline{x}),r^{\prime}(\cdot))$ is also Lipschitz continuous in $\mathcal{L}(F^{0})$. Let $L_{r}$ be the Lipschitz constant, meaning for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$
$\|u(\hat{x},\hat{\epsilon})-u(\tilde{x},\tilde{\epsilon})\|\leq L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$ (32)
A natural choice is $L_{r}=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$ and $L_{r}=r^{\prime\prime}(|\underline{x}|)$ if $r^{\prime}(|0^{+}|)=\infty$.
Next we prove that $S$ is Lipschitz continuous. It follows that
${\|}S^{x}(\hat{x},\hat{\epsilon})-S^{x}(\tilde{x},\tilde{\epsilon}){\|}=$ $\ {\|}\mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x})){\|}$
$\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\|\hat{u}-\tilde{u}\|$
$\leq$ $\ \|\hat{x}-\tilde{x}\|+\tfrac{1}{\beta}\|\nabla f(\hat{x})-\nabla f(\tilde{x}))\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f})\|\hat{x}-\tilde{x}\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
where the first inequality is determined by Lemma A.3 and the second is by (21) and (32). Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$. Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is invertible and a lipeomorphism. ∎
Now we are ready to apply Corollary 1 to derive the following global escape theorem.
Theorem 15.
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3 satisfied. Suppose $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$, $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$ and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
Proof.
It follows directly from Theorem 13 that the map $T$ is local $C^{1}$ smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points. By applying Proposition 14, $T$ is a lipeomorphism. By Corollary 1, the set of initial conditions attracted by such fixed points has a zero Lebesgue measure. ∎
We illustrate this property by applying DIRL1 to Example 1. We test $\alpha=0.1$ and $0.9$ with different initial points: $(3,3)$ representing an initial point far from the global minimizer and $(0,0.043)$ very close to the saddle point. The trajectories are shown in Figure 3. In four cases, the algorithm always converges to the global minimizer $(0,1)$.
(a) $\alpha=0.9$ and initial point $(3,3)$.
(b) $\alpha=0.9$ and initial point $(0,0.043)$.
(c) $\alpha=0.1$ and initial point $(3,3)$.
(d) $\alpha=0.1$ and initial point $(0,0.043)$.
Figure 3: Trajectories of DIRL1.
4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
The ${\rm IRL}_{2}$ algorithm is an effective method to solve problem (P). In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$. Since $\sqrt{\cdot}$ is concave on $\mathbb{R}_{+}$, $\sqrt{\lambda y_{1}+(1-\lambda)y_{2}}\geq\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}}$ for any $\lambda\in[0,1]$ and $y_{1},y_{2}\geq 0$. Since $r$ is increasing and concave, it follows that $h(\lambda y_{1}+(1-\lambda)y_{2})=r(\sqrt{\lambda y_{1}+(1-\lambda)y_{2}})\geq r(\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}})\geq\lambda r(\sqrt{y_{1}})+(1-\lambda)r(\sqrt{y_{2}})=\lambda h(y_{1})+(1-\lambda)h(y_{2})$, meaning $h$ is also concave on $\mathbb{R}_{+}$.
At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem (for brievty, we still use $S$ and $T$ for the maps),
$y^{k}=S^{x}(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2},$
where
$w(x_{i},\epsilon_{i})=h^{\prime}(x_{i}^{2}+\epsilon_{i}^{2})=\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{2\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}.$
In particular,
$w(x_{i},0)=\begin{cases}\frac{r^{\prime}(|x_{i}|)}{2|x_{i}|}&\text{ if }x_{i}\neq 0\\ \infty&\text{ if }x_{i}=0.\end{cases}$
The solution of the subproblem is given by
$y^{k}_{i}=S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i}^{k},\epsilon_{i}^{k})}(x_{i}^{k}-\frac{1}{\beta}\nabla_{i}f(x^{k})).$ (33)
If $w(x_{i}^{k},\epsilon_{i}^{k})=\infty$, we simply set $S_{i}^{x}(x^{k},\epsilon^{k})=0$. The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 10.
1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
2:Initialize: set $k=0$.
3:repeat
4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=h^{\prime}((x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})$.
5:     Compute new iterate:
6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2}.$
7:     $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$.
8:     $\epsilon^{k+1}=\mu\epsilon^{k}$.
9:     Set $k=k+1$.
10:until Convergence
Algorithm 2 Damped Iteratively Reweighted $\ell_{2}$ Algorithm (${\rm DIRL}_{2}$)
The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})$ with $T=(1-\alpha)I+\alpha S$ and
$T(x,\epsilon):=\begin{bmatrix}(1-\alpha)x+\alpha S^{x}(x,\epsilon)\\ (1-\alpha(1-\mu))\epsilon\end{bmatrix}.$ (34)
We have the following result to connect the stationary point of (P) and the fixed point of $T$.
Proposition 16.
If $x^{*}$ is a stationary point of (P), then $x^{*}=S^{x}(x^{*},0)$ and $(x^{*},0)=T(x^{*},0)$. Conversely, for any $x^{*}$ satisfying $x^{*}=S^{x}(x^{*},0)$, if $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, then $x^{*}$ is a stationary point of (P).
Proof.
Suppose $x^{*}$ is a stationary point. It follows from (6) that for $i\in\mathcal{I}^{*}$, $x_{i}^{*}=x_{i}^{*}-\tfrac{1}{\beta}(\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*}))$ implying $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$. Therefore, $x_{i}^{*}=S_{i}(x^{*},0)$ since $w(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$. As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$. In summary, $x^{*}=S^{x}(x^{*},0)$.
Now consider $x^{*}=S^{x}(x^{*},0)$. If $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, we only have to verify (6) is satisfied by $x^{*}$. It follows from $x^{*}_{i}=S_{i}(x^{*},0),i\in\mathcal{I}^{*}$ that $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$, implying $\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0$, completing the proof. ∎
The converse statement of Proposition 16 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$. Since for any $x^{*}$ that satisfies $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it is naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$. However, there is no guarantee that $x^{*}$ satisfies (7). In those cases, $S$ may have two types of fixed points — those are stationary points and others are not. This represents a stark difference from the DIRL1 algorithm, where all fixed points of the iteration map are stationary points.
4.1 Convergence analysis
In this subsection, we show that DIRL2 only converges to its fixed points which are the stationary points of (P) under the assumption that $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$. Similarly to the analysis of the DIRL1 algorithm, the analysis of DIRL2 is also based on the monotonic decrease in the objective $F(x,\epsilon)$. We also need Assumption 3 and consequently (21) to hold for the analysis. However, it should be noted that the definition of $F(x,\epsilon)$ is different from that of the DIRL1 algorithm.
Proposition 17.
Suppose Assumption 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
Proof.
It follows from the concavity of $h$ that for $i=1,...,n$,
$h(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2})\leq h(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})+h^{\prime}((x_{i}^{k})^{2}+(\epsilon^{k}_{i})^{2})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}),$
which is equivalent to
$r(\sqrt{(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2}})\leq r(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})+w(x_{i}^{k},\epsilon^{k}_{i})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}).$
Following the same argument for (24)-(26), we have
$\Delta F(x^{k+1},\epsilon^{k+1})\geq\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (35)
For each subproblem, $y^{k}$ satisfies the optimal condition
$\nabla f(x^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})=0.$ (36)
Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (27), the following holds
$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (37)
$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})((x_{i}^{k})^{2}-(x_{i}^{k+1})^{2})-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(\alpha(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2})-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+2\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ \alpha[\nabla f(y^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combining (26) and (37), we get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
Replacing $k$ with $t$ and summing up from $t=0$ to $k-1$, we have
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
It follows that $\{(x^{k},\epsilon^{k})\}\in\mathcal{L}(F^{0})$ since $F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$. By Assumption 3, it is bounded and $\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})>-\infty$. Taking $k\rightarrow\infty$, we have
$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$, there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
Theorem 18.
(Global convergence) Suppose Assumption 13 hold, and $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then any $x^{*}\in\Omega$ is first-order optimal for (P).
Proof.
Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})}{\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}}}y_{i}^{k}.$ (38)
As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$.
For $i\in\mathcal{J}^{*}$, there is nothing to prove since $|\nabla_{i}f(x^{*})|\leq r^{\prime}(|0^{+}|)$ naturally holds. In general, $x^{*}$ is stationary. ∎
4.2 Avoidance of strictly saddle points
In this subsection, we further show that the DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points. Unlike DIRL1, the map $S$ in DIRL2 does not enjoy the property of “active manifold identification” or “model identification”. There is no guarantee that the iteration map is $C^{1}$-smooth without this property. The following additional assumption is needed to guarantee the local smoothness. One can verify that most sparse nonconvex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
Assumption 4.
$\lim\limits_{z\rightarrow 0}r^{\prime}(|z|)=\infty\ \text{ and }\ \lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
To show the smoothness of $S$, we first calculate the Jacobian of $S$. Define
$g(z):=\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|z|)}{|z|}}\ \text{ with }g(0):=0.$
We can write $S_{i}^{x}(x,\epsilon)=g(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ with $z=\sqrt{x^{2}+\epsilon^{2}}$. It follows that
$g^{\prime}(z)=\begin{cases}-\tfrac{\lambda}{\beta}\frac{r^{\prime\prime}(|z|)z-\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}},&\text{\ for }z\neq 0,\\ \lim\limits_{\delta\to 0}\left(\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|\delta|)}{|\delta|}}-0\right)\frac{1}{\delta}=0,&\text{\ for }z=0.\end{cases}$ (39)
It holds that
$\lim\limits_{z\to 0}g^{\prime}(z)=$ $-\tfrac{\lambda}{\beta}\lim\limits_{z\to 0}\left(\frac{r^{\prime\prime}(|z|)z}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}-\frac{\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}\right)$
$=$ $-\lim\limits_{z\to 0}\left(\tfrac{\beta}{\lambda}\frac{r^{\prime\prime}(|z|)z}{{r^{\prime}(|z|)}^{2}}-\text{sign}(z)\frac{|z|}{r^{\prime}(|z|)}\right)=0=g^{\prime}(0)$
by Assumption 4. Therefore, $g$ is continuously differentiable in $\mathbb{R}$.
We have the following result for $S$ and $T$.
Theorem 19.
(Local smoothness and Lipschitz continuity) Suppose Assumption 1-4 hold. The following statements hold:
  1. (i)
    $S$ is $C^{1}$ smooth and Lipschitz continuous on $\mathcal{L}(F^{0})$ with constant $L_{S}$; moreover
    $DS(x^{*},\bm{0})=\ \begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&,\bm{0}\\ 0&0&0\\ 0&0&\mu I\end{bmatrix},$ (40)
    where $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$.
  2. (ii)
    $T$ is a lipeomorphism for $\alpha\in(0,\frac{1}{1+L_{S}})$ with
    $DT(x^{*},\bm{0})=\ \begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}.$
Proof.
(i) We simply calculate its partial derivative and verify that it is continuous at any point $(x^{*},\epsilon^{*})\in\mathbb{R}^{2n}$.
Consider $i$ satisfying $z_{i}^{*}:=\sqrt{(x_{i}^{*})^{2}+(\epsilon_{i}^{*})^{2}}>0$. In this case, $S$ is smooth with respect to $x_{i}$ and $\epsilon_{i}$, and
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ g(z^{*}_{i})(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))+\tfrac{x_{i}^{*}}{z_{i}^{*}}g^{\prime}(z_{i}^{*})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (41)
$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ -g(z^{*}_{i})\tfrac{1}{\beta}\nabla^{2}_{ij}f(x^{*}),\quad j\neq i;$ (42)
$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ \tfrac{\epsilon_{i}^{*}}{z_{i}^{*}}g^{\prime}(z^{*}_{i})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (43)
$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ 0,\quad j\neq i.$ (44)
Consider $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$. It follows from $S_{i}^{x}(x^{*},\bm{0})=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$ (45)
$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta$
$=$ $-\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}g^{\prime}(0)=0,$
$\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{j})]/\delta$ (46)
$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0,\quad j\neq i;$
$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(0,\delta)}[-\tfrac{1}{\beta}\nabla_{i}f(x^{*})]/\delta$ (47)
$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0;$
[…]
Now we verify that the Jacobian is continuous at $(x^{*},\epsilon^{*})$. It suffices to show this for any $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$; indeed, it holds that It follows from (41), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g(0)=0,g^{\prime}(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}).$
It follows from (42) and $g(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*}).$
It follows from (43), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g^{\prime}(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
It follows from (44) that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
Overall, we have shown that $S$ is continuously differentiable.
To derive the matrix form of the Jacobian, it follows from (41) that for $i\in\mathcal{I}^{*}$,
$h_{ii}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})$
$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(x^{*}_{i}))\tfrac{\lambda}{\beta}\left(|x^{*}_{i}|r^{\prime\prime}(|x^{*}_{i}|)-r^{\prime}(|x^{*}_{i}|)\right)}{\left(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(|x^{*}_{i}|)\right)^{2}},$
$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x_{i}^{*}|}}$
$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x_{i}^{*}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x_{i}^{*}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
It follows from (42) that for $j\neq i$,
$h_{ij}:=\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
We then combine this with (43)–(48) to derive (40) of the Jacobian $DS(x^{*},\bm{0})$. The Lipschitz continuity of $S$ trivially leads to the boundedness of $\mathcal{L}(F^{0})$.
(ii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. The Jacobian $DT(x^{*},\bm{0})$ is obtained directly by (i). ∎
Now we show that every strict saddle point of (P) is an unstable fixed point of $T$.
Theorem 20.
(Unstable Fixed Point) Suppose Assumption 1-4 hold. Let $x^{*}$ be a stationary point of (P). If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=\bm{0}$ is an unstable fixed point of $T$. Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible.
Proof.
It follows from Theorem 19(ii) that we only have to check the eigenvalues of
$(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}),$ (49)
where we have $W^{*}_{ii}=w(x_{i}^{*},0)<\infty,i\in\mathcal{I}^{*}$. If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [49, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$. Hence, $DT(x^{*},\bm{0})$ has an eigenvalue greater than 1, meaning $(x^{*},\bm{0})$ is the unstable fixed point of $T$. Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible. ∎
Now we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
Theorem 21.
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{2}$ algorithm with Assumption 1-4. Let the condition of Proposition 19 holds and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
Proof.
It follows directly from Theorem 19 and 20 that: (i) map $T$ is local $C^{1}$-smooth around any stationary points; (ii) all stationary points $x^{*}$ are unstable fixed points; (iii) $T$ is a lipeomorphism. Then by Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure. ∎
We illustrate this property by applying DIRL2 to Example 1. We test $\alpha=0.1$ and $0.9$ with different initial points; first is $(3,3)$ representing an initial point far from the global minimizer and the second is $(0,0.043)$ very close to the saddle point. The trajectories are shown in Figure 4. In four cases, the algorithm always converges to the global minimizer $(0,1)$.
(a) $\alpha=0.9$ and initial point $(3,3)$.
(b) $\alpha=0.9$ and initial point $(0,0.043)$.
(c) $\alpha=0.1$ and initial point $(3,3)$.
(d) $\alpha=0.1$ and initial point $(0,0.043)$.
Figure 4: Trajectories of DIRL2.
Appendix
Lemma A.1.
Given $a,b,c\in\mathbb{R}^{n}_{+}$ with $a_{i}\leq b_{i}+c_{i}\quad\text{for all }i$. It holds that $\|a\|\leq\|b\|+\|c\|$.
Proof.
It follows that $a_{i}^{2}\leq b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2}$, implying $\sum_{i=1}^{n}a_{i}^{2}\leq\sum_{i=1}^{n}(b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2})$, or equivalently, $\|a\|^{2}\leq\|b\|^{2}+2b^{T}c+\|c\|^{2}$. By the Cauchy-Schwarz inequality, $2\sum_{i=1}^{n}b_{i}c_{i}\leq 2\|b\|\|c\|$. Therefore, $\|a\|^{2}\leq(\|b\|+\|c\|)^{2}$, completing the proof. ∎
Lemma A.2.
For $x,y\in\mathbb{R}$ and $\alpha\in(0,1]$, it holds that $x^{2}-((1-\alpha)x+\alpha y)^{2}\geq\alpha(x^{2}-y^{2})$.
Proof.
It follows that
$x^{2}-((1-\alpha)x+\alpha y)^{2}$ $=x^{2}-(1-\alpha)^{2}x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
$=(2\alpha-\alpha^{2})x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})x^{2}+(\alpha-\alpha^{2})y^{2}-2(\alpha-\alpha^{2})xy$
$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})(x-y)^{2}$
$\geq\alpha(x^{2}-y^{2}),$
where the inequality is by $\alpha\in(0,1]$. ∎
Lemma A.3.
Given $\hat{w},\tilde{w}\in\mathbb{R}^{n}_{+}$ and $\hat{z},\tilde{z}\in\mathbb{R}^{n}$, it holds that
$\|\mathbb{S}_{\hat{w}}(\hat{z})-\mathbb{S}_{\tilde{w}}(\tilde{z})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$
Proof.
(a) If $|\hat{z}_{i}|\leq\hat{w}_{i}$ and $|\tilde{y}_{i}|\leq\tilde{w}_{i}$, then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=0$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$. In this case, $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=|0-0|=0\leq|\hat{z}_{i}-\tilde{z}_{i}|$.
(b) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|\leq\tilde{w}_{i}$ (or vice versa), then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$. In this case, $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=||\hat{z}_{i}|-\hat{w}_{i}|$. It follows that
$|\hat{z}_{i}-\tilde{z}_{i}|\geq$ $\ |\hat{z}_{i}|-|\tilde{z}_{i}|\geq|\hat{z}_{i}|-\tilde{w}_{i}$
$=$ $\ ||\hat{z}_{i}|-\hat{w}_{i}+\lambda(\hat{w}_{i}-\tilde{w}_{i})/\beta|$
$\geq$ $\ |\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|-|\lambda(\hat{w}_{i}-\tilde{w}_{i})/\beta|,$
implying
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
(c) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|>\tilde{w}_{i}$, then
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})-\text{sign}(\tilde{z}_{i})(|\tilde{z}_{i}|-\tilde{w}_{i})|.$
If $\text{sign}(\hat{z}_{i})=\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|(|\hat{z}_{i}|-\hat{w}_{i})-(|\tilde{z}_{i}|-\tilde{w}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|$. If $\text{sign}(\hat{z}_{i})=-\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=(|\hat{z}_{i}|-\hat{w}_{i})+(|\tilde{z}_{i}|-\tilde{w}_{i})\leq|\hat{z}_{i}|+|\tilde{z}_{i}|=|\hat{z}_{i}-\tilde{z}_{i}|.$
In all three cases, we show
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
Therefore, by (A.1),
$\|\mathbb{S}_{\hat{w}}(\hat{w})-\mathbb{S}_{\tilde{w}}(\tilde{w})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$